### v1

# Predictive Reliability Assessment of Distribution Grids with Residential Distributed Energy Resources

Abstract
Distribution system end users are transforming from passive to active participants, marked by the push towards widespread adoption of edge-level Distributed Energy Resources (DERs). This paper addresses the challenges in distribution system planning arising from these dynamic changes. We introduce a bottom-up probabilistic approach that integrates these edge-level DERs into the reliability evaluation process. Our methodology leverages joint probability distributions to characterize and model the penetration of rooftop photovoltaic (PV) systems and energy storage across a distribution network at the individual residential level. Employing a scenario-based approach, we showcase the application of our probabilistic method using a Monte Carlo Simulation process to assess average system reliability indices and their variations at the user level. To validate our approach, we applied this methodology to the RBTS test system across various adoption scenarios, effectively showcasing the capability of our proposed method in quantifying the variation in end-user reliability indices for each scenario within the distribution system.
Index Terms:
Distributed energy resources, predictive reliability assessment, distribution grids.
I Introduction
Distribution systems have transitioned from unidirectional, static, radial systems [1] to bi-directional, dynamic, and meshed systems and evolved to become complex interconnected systems. The importance of distribution system reliability for utilities/distribution system operators (DSO) cannot be overstated, encompassing diverse areas from community welfare to asset management. Beyond their primary function of transporting electricity, distribution systems are foundational in upholding public health and fostering a community’s economic vitality [2]. Consequently, maintaining their efficient performance is not a mere operational necessity but is crucial for the broader well-being of communities. Utilities increasingly apply asset management techniques to their distribution systems to identify factors influencing infrastructure reliability to make informed decisions about their systems’ maintenance, upgrades, and investments [3].
Customer/end user satisfaction, an integral measure for utilities, is directly influenced by system reliability. Utilities ensure a consistent power supply by assessing reliability and fostering an environment conducive to economic development and heightened customer satisfaction. Furthermore, reliability assessments help utilities optimize investments in distribution systems, minimizing the costs of interruptions to both the utility and its customers, leading to a more cost-effective and efficient distribution system [4].
There are two types of reliability assessments: historical and predictive. The objective of historical assessment is to measure and report the system’s actual performance, identifying areas of improvement based on past and current data. On the other hand, predictive reliability assessment aims to quantify potential reliability issues based on projections, scenarios, and anticipated changes in the system. Its objective is to guide the expansion and development of the distribution system in planning and designing systems that can adapt to future changes and challenges. This paper is motivated by the need for a predictive assessment method that integrates the end user DER into the reliability assessment.
I-A Related Works
As DERs became an emerging technology in distribution systems, the focus was on integrating them at the system/feeder level. In this context of reliability assessment, the research focused on integrating these resources into the assessment framework [5]. Both analytical [6, 7, 8, 9, 10] and simulation methods [11, 12, 13, 14, 15, 16, 17], were developed to align with the existing literature. For instance, [6] presents a generalized analytical formulation for reliability evaluation with DGs and microgrids, and [7] proposes an encoded Markov set algorithm for reliability evaluation with both conventional and renewable DG. [8] proposes an analytical method considering the protection strategy, islanding operation, and supply restoration process in the presence of DGs.
Similarly, sequential [11, 12] and non-sequential simulation [15] methods were developed to integrate Distributed energy resources (DERs) into reliability evaluation, considering various aspects such as energy storage integration [11], the islanding dynamics of microgrids [15], and electric vehicle integration [13, 16, 17]. All these approaches evaluated the primary load point reliability indices, namely failure rate ($\lambda$), repair time ($r$), and unavailability ($U$) in the presence of DGs; these load point indices were used to calculate the system reliability indices using the formulation in [18]. [19], present a three level hierarchical reliability assessment framework to integrate DERs both at the system/feeder level and the end user level.
Limitations of the current reliability assessment methods: Behind-the-meter (BTM) DERs allow end users to play a dynamic role as a sink and source based on the operating conditions of the distribution system. Given this new paradigm in the distribution system, using primary load point indices to evaluate reliability is myopic. The following example elucidates the limitations of the current approach. Consider a system with two load points and five residential customers connected to each of them. Load point data is provided in Table I.
TABLE I: Load point data for Example system
Load points customers $\lambda$ (f/y) U (hrs)
LP1 [1,2,3,4,5] 3 5
LP2 [6,7,8,9,10] 2 10
The reliability indices for the example system evaluated by the current approach of using the basic load point indices are as follows:
$\text{SAIFI}_{\text{P}}=\frac{\sum\lambda_{i}C_{i}}{\sum C_{i}},\quad\text{SAIDI}_{\text{P}}=\frac{\sum U_{i}C_{i}}{\sum C_{i}}$ (1)
The subscript ’P’ here means ’perceived’ indicating that the reliability metrics evaluated are as seen by the DSO/utility side. The actual customer experience can be quantified using the same reliability indices as follows:
$\text{SAIFI}_{\text{E}}=\frac{\sum_{C}\text{Interruptions experienced }}{\sum C}$ (2)
$\text{SAIDI}_{\text{E}}=\frac{\sum_{C}\text{Interruption durationexperienced }}{\sum C}$ (3)
The subscript ’E’ means experienced indicating that the reliability metrics evaluated are as seen from the customer side of the distribution system. Consider the following two cases for the example system:
Figure 1: Case1: NO DER, Case2: DER at 4,5,9,10.
Case 1: There are no BTM DERs in this case. The failure rate and the unavailability of the load point are equal to the interruption frequency and the restoration time of the end user, respectively. Therefore, the system indices evaluated using the load point approach and those evaluated from customer experience yield the same numerical result. Similar results indicate that DSO’s perceived reliability is same as the reliability experienced by the customers, as shown in Fig. 1.
Case 2: 40% of the customers have BTM DER (customers 4,5,9 and 10). As shown in Fig. 1, for customers 4,5,9 and 10, BTM DER can meet the system demand whenever there is a load point failure, therefore they never experience an interruption. Given this context, the system indices evaluated using the load point approach by the DSO and that evaluated from customer experience yield different numerical results. The disparity in the indices indicates that the reliability perceived by the DSO is not the same as the reliability experienced by the customers. While the example assumes simplistic scenarios, it shows that the current approach of using the basic load point indices to evaluate system reliability does not correctly quantify the reliability value provided by BTM DER assets.
I-B Motivation and Scope of this Work
In the evolving landscape of modern distribution systems, the proliferation of BTM DERs challenges traditional notions of reliability evaluation. Modern energy consumers, driven by rapidly shifting paradigms, demand more than just power supply. Their expectations encompass flexibility, resilience, and active participation in the energy market through DERs. This active involvement can significantly influence customer satisfaction levels, bridging or widening the gap between utility offerings and end-user expectations. Relying on traditional financial evaluations might lead utilities down paths that, while appearing cost-effective at first glance, do not account for the full spectrum of value offered by DERs, potentially resulting in subpar investment decisions. The dynamic integration of DERs affects wear and tear, performance, and asset management strategies. By overlooking their influence, utilities may inadvertently miss signals crucial for timely maintenance, putting the system’s integrity and reliability at risk.
Furthermore, the contemporary energy transition, driven in part by the integration of DERs, offers an unprecedented opportunity to address longstanding disparities in energy access and costs. Disregarding these resources risks perpetuating past inequities, a misstep starkly contrasting the ideals of a modern, inclusive grid. As the role of end-users undergoes a transformation from passive consumption to active production and integration, reliability assessments must capture this significant shift.
This paper builds on the existing literature on distribution system reliability assessment. The proposed method represents a modular addition to the existing assessment methods, designed to integrate seamlessly and augment their functionality. In this bottom-up approach, we first provide a residential system model that can be configured based on the BTM DER. We provide the formulations for the reliability evaluations for the residential system. We developed the probabilistic approach and the required DER penetration modeling techniques for predictive reliability assessment. We applied the proposed method to a modified RBTS test system and critically analyzed the results. This analysis demonstrated the practical application of the approach, revealing its effectiveness in various scenarios. The results underscore the utility of the probabilistic method in enhancing the accuracy and comprehensiveness of reliability assessments in modern distribution systems. They also highlight the method’s ability to account for the complexities and variabilities introduced by DERs, showcasing its potential as a valuable tool for future-oriented energy system planning and management. The contributions of this paper are summarized as follows:
  • We propose a bottom-up probabilistic approach to evaluate the reliability of distribution systems using a general-purpose modular residential system configuration.
  • We define a novel way of modeling DER penetration utilizing probability distributions to model micro-level penetration at the granularity of the residential level.
  • We propose a Monte Carlo Simulation method to estimate distribution system reliability indices and their underlying distributions at the residential level.
The remainder of the manuscript is organized as follows: Section 2 describes the residential system reliability evaluation. Section 3 proposes the probabilistic approach to system reliability evaluation. Section 4 introduces the DER penetration model. Section 5 describes the simulation method to evaluate system reliability and application of the methodology to RBTS test system.
II Residential system Reliability evaluation
To estimate the reliability of a distribution system with residential DER, it is imperative to understand the reliability of the residential system and the factors affecting it. Quantifiable metrics for the residential system reliability evaluation are Average interruption frequency (AIF), Average interruption duration (AID), and Average energy not served (AENS).
II-A Residential system model
In this paper, we define a general purpose configuration of a residential system. This configuration has three main components: the load, the DER system, and the Energy management system (EMS). The EMS controls the DER operations and connection to the grid. The EMS operates the residential system primarily in self-consumption mode, where the DER system acts as the primary supply and the utility connection acts as a backup to meet demand not served by the DER. The popular choice of DERs for residential systems is the rooftop PV and ES, and the general purpose configuration reflects both rooftop PV and ES, as shown in the figure 2. The general purpose model can be reduced to specific configurations based on the DER employed, such as PV only, ES only, or a no DER case where the configuration reduces to a regular residential system.
II-A1 PV system model
The output of the PV system at any time-step $t$ is modeled as follows,
$\mathrm{PV}_{\text{out}}(t)=\mathrm{PV}_{\text{cap}}\cdot d\cdot\mathrm{GHI}(t)$ (4)
where, $\mathrm{PV}_{\text{out}}$ is the output from the rooftop PV system, $\mathrm{PV}_{\text{cap}}$ is the nameplate capacity, $d$ is the derating factor and $\mathrm{GHI}$ is the global horizontal irradiation at the location of the PV system.
Figure 2: General-purpose residential system configuration.
II-A2 ES system model
Energy storage is modeled using the State of Charge (SOC) of the system where SOC is percentage that represents the fraction of the maximum capacity.
$\mathrm{SOC}(t)=\mathrm{SOC}(t-1)+\frac{\left(\eta_{c}Ch(t)-\frac{D(t)}{\eta_{d}}\right)}{\mathrm{ES}_{\text{cap}}}$ (5)
$\mathrm{ES}(t)=\mathrm{SOC}(t)\cdot\mathrm{ES}_{\text{cap}}$ (6)
$0\leq Ch(t)\leq Ch_{\text{max}}$ (7)
$0\leq D(t)\leq D_{\text{max}}$ (8)
$0\leq\mathrm{SOC}(t)\leq\mathrm{SOC}_{\text{max}}$ (9)
where, $\eta_{c}$ and $\eta_{d}$ are the charging and discharging efficiency respectively, $Ch$ and $D$ are the charging and discharging energy at a given time-step and $\mathrm{ES}$ is the available energy within the storage at a given time-step.
II-A3 Residential system energy balance
The energy balance for the residential system is modeled as follows:
$l_{n}(t)=l(t)-\mathrm{PV}_{\text{out}}(t)\pm\mathrm{ES}(t)$ (10)
where, $l(t)$, represents the load of the residential system and $l_{n}(t)$ represents the net load of the residential system at time $t$. Net load is the load seen by the utility at the residential system utility meter.
II-B Residential system reliability metrics
In the context of the residential system model presented, these reliability indies are evaluated as follows: Net load determines the state of the energy balance within the residential system and it has three states. Excess energy: When the PV produces more energy than the residential system can consume or store. Unserved load: When the PV output and the ES combined cannot meet the residential system consumption. Balanced load: When the PV output and the ES system operation perfectly meet the residential system consumption.
To put this into perspective, form the energy balance equation in (10), if $l_{n}(t)$ is positive, it means there is an unserved load and energy is drawn from the distribution system. If $l_{n}(t)$ is negative, it indicates the availability of excess energy which can be fed back to the distribution system or curtailed, and zero indicates that the onsite energy generation meets the load.
We define a binary state variable to represent the state of the residential system energy balance as follows:
$s_{\text{n}}(t)=\begin{cases}0&l_{n}(t)>0\\ 1&l_{n}(t)\leq 0\end{cases}$ (11)
The reliability metrics not just depend on the state of the residential system energy balance but also depend on the state of the load point to which the residential system is connected. The state of the load point depends on the reliability characteristics of the load point. The load point reliability characteristics are usually defined by the load point reliability metrics evaluated using the current methods and the synthetic sequential load point history for a given time horizon can be generated by sequentially calculating the up and down times of the load point as follows:
$\mathrm{T}_{\text{TTF}}=-\frac{\ln(\mathrm{R}_{1})}{\lambda_{\text{lp}}}$ (12)
$\mathrm{T}_{\text{TTR}}=-\frac{\ln(\mathrm{R}_{2})}{\mu_{\text{lp}}}$ (13)
where, $\mathrm{T}_{\text{TTF}}$ is the time to failure of the load point, $\mathrm{T}_{\text{TTR}}$ is the time to repair of the load point, $\lambda_{\text{lp}}$ is the failure rate, and $\mu_{\text{lp}}$ is the repair rate of the load point and $\mathrm{R}$ is a random number drawn from a uniform distribution . We define a binary state variable to describe the state of the load point at time t as follows:
$s_{\text{lp}}(t)=\begin{cases}0&\text{if LP is down}\\ 1&\text{if LP is up}\end{cases}$ (14)
Given the state of the energy balance and the state of the load point, the residential system state can be defined as follows. Operating State: When either the DER meets the local load or the load point is in the operating condition. Failed State: When there is unserved load and the load point is in the failed condition. We define another state variable to describe the state of the residential system as follows:
$s_{r}=s_{lp}\lor s_{n}$ (15)
Given the residential system state variable $s_{r}$, an interruption event occurs when there is a transition from the operating state ($s_{r}(t-1)=1$) to the failed state ($s_{r}(t)=0$). The logical expression capturing this transition is given as follows:
$\mathrm{I}(t)=s_{r}(t-1)\land\neg s_{r}(t)$ (16)
where $\mathrm{I}$ is the Boolean variable representing the interruption event at time $t$. This variable is True for every time $t$, where an interruption starts. The AIF is then calculated by averaging the number of interruption events over the study period as follows:
$\mathbf{AIF}=\frac{\sum_{t=1}^{T}\mathrm{I}}{T}$ (17)
The AID is calculated by averaging the duration of interruption over the study period as follows:
$\mathbf{AID}=\frac{\sum_{t=1}^{T}s_{r}}{T}$ (18)
We used a Monte Carlo simulation method to estimate the reliability indices of the residential system [20].
III Probabilistic approach to system reliability assessment
This section provides a detailed description of the bottom up probabilistic approach to evaluate the reliability of the distribution system using the general purpose residential system configuration as a basic building block. In the context of system reliability, the distribution system is viewed as a set of load points and the associated customers connected to the load point.
We define the set of load points as $\mathcal{L}$, where
=$\mathcal{L}=\{L_{1},L_{2},\dots,L_{n}\}$ (19)
Each $L_{i}$ represents a specific load point for $i=1,2,\dots,n$. For each load point $L_{i}$, there is an associated set of customers $C_{i}$ connected to it.
$C_{i}=\{c_{i1},c_{i2},\dots,c_{im_{i}}\}$ (20)
where $m_{i}$ is the number of customers connected to the load point $L_{i}$ and each $c_{ij}$ represents an individual customer connected to $L_{i}$. To represent all the sets of customers connected to the load points, we define
$\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$ (21)
where $C_{i}$ is the set of customers connected to the corresponding load point $L_{i}$. The relation ship between the set of load points and the set of customers is captured as follows:
$\forall L_{i}\in\mathcal{L},\exists C_{i}\in\mathcal{C}$ (22)
Residential system reliability metrics quantify the susceptibility to interruptions and are influenced by several parameters, namely $l$, $\mathrm{PV}_{cap}$, $\mathrm{ES}_{cap}$, and the reliability of the load point which is quantified by the state of the load point $s_{\text{lp}}$. Together, these variables articulate the circumstances under which the residential system may experience failures, thereby contributing to the residential system reliability metrics. Using the defined notation, we can express the reliability metrics as a function of the influencing variables as follows.
$\mathbf{AIF}_{c_{ij}}=g\left(\mathrm{PV}_{c_{ij},\text{cap}},\;\mathrm{ES}_{c_{ij},\text{cap}},\;l_{c_{ij}},\;s_{L_{i},\text{lp}}\right)$ (23)
$\mathbf{AID}_{c_{ij}}=h\left(\mathrm{PV}_{c_{ij},\text{cap}},\;\mathrm{ES}_{c_{ij},\text{cap}},\;l_{c_{ij}},\;s_{L_{i},\text{lp}}\right)$ (24)
$\forall c_{ij}\in\mathcal{C}$
For the purpose of illustration, only $\mathbf{AIF}$, is used to show the bottom up formulation to evaluate system metrics. At this point, we introduce different layers of abstraction based on the granularity of the analysis.
Abstraction layer 1: In the first layer, the reliability provided by the load points is uniform across the distribution system. In the context of predictive reliability assessment, it is a reasonably fair abstraction given the goal is to quantify the BTM residential DER contribution to the distribution system reliability. Therefore, $s_{lp}$ has a consistent influence on the $\mathbf{AIF}_{c_{ij}}$, allowing us to simplify the function to:
$\mathbf{AIF}_{c_{ij}}=g\left(\mathrm{PV}_{c_{ij},\text{cap}},\mathrm{ES}_{c_{ij},\text{cap}},l_{c_{ij}}\right)$ (25)
To address the disparity in load magnitude, PV system sizes, and energy storage capacities among different residential systems, we normalize these variables with the peak load ($l_{c,p}$), yielding:
$\begin{split}\mathbf{AIF}_{c_{ij}}&=g\left(\frac{\mathrm{PV}_{c_{ij},\text{cap}}}{l_{c_{ij},p}},\;\frac{\mathrm{ES}_{c_{ij},\text{cap}}}{l_{c_{ij},p}},\;\frac{l_{c_{ij}}}{l_{c_{ij},p}}\right)\\ &=g\left(\mathrm{PV^{\prime}},\;\mathrm{ES^{\prime}},\;l^{\prime}\right)\end{split}$ (26)
The above equation means that for any residential unit connected to the distribution system, the $\mathbf{AIF}$ is a function of the load profile $l^{\prime}$ and the normalized PV and ES system sizes. Assuming homogeneity in load profiles across residential systems, with variations in load values being primarily due to the magnitude of the load, $l^{\prime}$ is consistent for all residential systems, leading to a further simplified model:
$\mathbf{AIF}_{c_{ij}}=g\left(\mathrm{PV^{\prime}},\;\mathrm{ES^{\prime}}\right)=g\left(X,\;Y\right)$ (27)
for simplicity we define $\mathrm{PV^{\prime}}=X$ and $\mathrm{ES^{\prime}}=Y$, where $X$ and $Y$ as continuous random variables with a joint density $f_{XY}(x,y)$, representing the penetration or forecasted penetration of DERs among the residential systems connected in the distribution system. The SAIFI is simply the expected value of the $\mathbf{AIF}$ over the joint density of $f_{XY}(x,y)$
$\begin{split}\mathbf{SAIFI}&=\mathop{\mathbb{E}}\left(\mathbf{AIF}_{c_{ij}}\right)\\ &=\mathbb{E}\left(g(X,\;Y)\right)\end{split}$ (28)
where, the expectation of $g(X,Y)$ with respect to the joint density is given by
$\mathbb{E}\left[g(X,Y)\right]=\iint g(x,y)f_{XY}(x,y)\,\mathrm{d}x\,\mathrm{d}y$ (29)
Abstraction layer 2: In the second layer of abstraction, the reliability provided by the load points is non uniform across the distribution system. Therefore (23) is simplified to the following:
$\mathbf{AIF}_{L_{i}}=g\left(\mathrm{PV}_{c_{ij},\text{cap}},\mathrm{ES}_{c_{ij},\text{cap}},l_{c_{ij}}\right)$ (30)
$\forall c_{ij}\in C_{i}$
where, $s_{lp}$ has a consistent influence on the residential systems connected to the load point $L_{i}$. Further simplification as in the earlier abstraction layer will lead to
$\mathbf{AIF}_{L_{i}}=g\left(\mathrm{PV^{\prime}},\;\mathrm{ES^{\prime}}\right)=g\left(X,\;Y\right)$ (31)
in this case the joint density $f_{L_{i},{XY}}(x,y)$ represents the penetration or forecasted penetration of DERs across the residential systems connected to the load point $L_{i}$.
We define the load point average interruption frequency index for a load point $L_{i}$ as the expected value of $\mathbf{AIF}$ over the residential systems connected to $L_{i}$. It simply is the SAIFI for the load point $L_{i}$
$\begin{split}\mathbf{LAIFI}_{L_{i}}&=\mathop{\mathbb{E}}\left(\mathbf{AIF}_{L_{i}}\right)\\ &=\iint g(x,y)f_{L_{i},{XY}}(x,y)\,\mathrm{d}x\,\mathrm{d}y\end{split}$ (32)
The model for the load point is then extended to a distribution system with multiple load points each with a different expectation of the AIF. System Average Interruption Frequency Index is obtained by taking a weighted average of each load point’s average interruption frequency.
$\mathbf{SAIFI}=\sum_{i=1}^{n}w_{L_{i}}*\mathbf{LAIFI}_{L_{i}}$ (33)
$w_{L_{i}}=\frac{|{C_{i}}|}{\sum_{i=1}^{n}|C_{i}|}$ (34)
Similar to these two abstraction layers, another layer that could be of practical importance is at the feeder level. In this layer, the reliability provided by the load points is uniform across a distribution feeder.
In all the layers of abstraction, we are primarily quantifying the reliability using an average metric such as $\mathbf{SAIFI}$. While the average provides an excellent basis to quantify the contribution of DERs to system reliability, it would be more insightful and valuable to understand the underlying distribution of the residential reliability metric for a given distribution of DER penetration. From (27) we can see that the random variable $\mathbf{AIF}$ is a transformation of the random variables $X$ and $Y$ through the function $g$. To determine the probability distribution of $\mathbf{AIF}$, we can leverage the joint PDF $f_{XY}(x,y)$. Since an explicit form of $g$ is not available, numerical methods such as Monte Carlo Simulations combined with kernel density estimation can be employed to approximate the desired PDF of the reliability metric.
IV Modeling DER penetration
In the proposed probabilistic model, the necessary information to evaluate system reliability is the joint density $f_{XY}(x,y)$. This density refers to the penetration of edge level DERs at the residential level. The advantages of such modeling, and the technical details are presented in the following subsections.
IV-A Macro vs Micro Perspectives
DER penetration is usually quantified with aggregates at the bulk system level or stratified into different customer sectors at the distribution system level. It is defined as the total installed capacity of the DER across a system, relative to the peak load of that system.
In the context of PV and ES, they are defined as follows:
$\phi_{\text{PV}}=\frac{\text{Total Installed PV capacity (kW)}}{\text{Total peak load(kW)}}$ (35)
$\phi_{\text{ES}}=\frac{\text{Total Installed ES capacity (kWh)}}{\text{Total peak load(kW)}}$ (36)
These macro penetration metrics $\phi_{\text{PV}}$, $\phi_{\text{ES}}$ provide a holistic view of how saturated a given system or region is with the adopted DER technology. While useful, this high-level perspective can obscure variations and patterns at the residential levels. In this rapidly evolving landscape of edge-level DERs, understanding the penetration of DER can no longer remain a macro affair. Instead, capturing the intricacies of individual behaviors, choices, and motivations is paramount as the actual dynamics of energy transition play out at this granular edge level. By modeling the penetration at the micro-level, one can peel back the layers of aggregate statistics to reveal the vibrant mosaic of individual decisions and factors that drive the bigger picture.
IV-B Quantitative Modeling of Micro-level Penetration
The primary goal is to model the micro-level penetration of DER installations at individual residential units within a distribution system. By characterizing this penetration as a probabilistic phenomenon, we aim to make informed predictions, simulate scenarios, and derive insights about the distribution system’s reliability.
Let $X$ and $Y$ represent the PV and ES penetration of a randomly chosen residential system from a defined distribution system, where
$X=\frac{\mathrm{PV}_{\text{cap}}}{l_{\text{p}}}$ (37)
$Y=\frac{\mathrm{ES}_{\text{cap}}}{l_{\text{p}}}$ (38)
$X$ indicates how the rooftop PV system capacity compares to the residential system peak electricity demand.
  • $X=1$: PV system capacity matches the peak load
  • $X>1$: PV system has excess capacity relative to the peak load
  • $X<1$: PV system might not fully cover the peak load
$Y$ indicates how much energy the ES system can hold relative to the residential system peak electricity demand.
  • $Y=1$: ES system can support the residential system peak load for an hour.
  • $Y>1$: ES system can support the residential system peak load for more than an hour.
  • $Y<1$: ES system cannot supply the residential system peak load for a full hour.
Together, these ratios provide insight into how self-sufficient a residential system might be in terms of generating and storing its own electricity.
We model the random variables X and Y as probability distributions to characterize the variations and patterns in DER installations among distinct residential units within a distribution system. The probability distribution parameters will describe specific aspects of the micro-level DER penetration within the distribution system.
Examining the marginal distributions of $X$ and $Y$ yields insights into individual characteristics; however, the intricate dynamics and the co-dependence between rooftop PV and ES systems necessitate joint modeling to understand their combined implications on the distribution system. In this work, we construct scenarios for the joint adoption of $X$ and $Y$ and assign specific correlations, allowing us to explore various ways these variables might interact.
V Application of the probabilistic approach
In this section, we apply the probabilistic method to analyze the RBTS test system. We begin by thoroughly explaining the Monte Carlo Simulation framework used to determine the reliability indices and their distributions. Following that, we detail the marginal adoption distributions for PV and ES and the modeling of their joint scenarios. We conclude the section with a presentation and critical analysis of the simulation results.
V-A MCS Simulation
We propose a Monte Carlo simulation (MCS) framework that leverages the probabilistic model discussed earlier to estimate the distribution of reliability indices such as AIF and AID across the distribution system. The following steps outline the MCS:
  1. 1.
    Sample Generation: Draw samples from the joint probability distribution of X and Y.
  2. 2.
    Index calculation: Each sample corresponds to a specific configuration of X and Y for the residential system. Evaluate the desired reliability metric for every unique system configuration derived from the samples.
  3. 3.
    Stopping criterion: Employ the confidence interval width as the stopping criterion. Continue simulations until the width of the CI is within a predefined precision threshold.
The detailed simulation flowchart is shown in Fig. 3. The steps described above are the generalized MCS steps for the abstraction where the reliability characteristics of the load points are uniform across the distribution system. In the case of more granular abstraction layers, such as the one where each load point has different reliability characteristics, the same set of steps detailed earlier are used to evaluate the index at each load point, and the system index is then calculated using the weighted average of the load point indices as expressed in (33).
Figure 3: MCS flowchart.
V-B Marginal Distributions for X and Y
Field penetration data provides a solid foundation for accurately modeling DER penetration. However, a significant challenge arises due to the shallow adoption of BTM DER, especially in energy storage in the current energy landscape. This shallow adoption hinders our ability to construct meaningful, data-driven adoption scenarios. Consequently, we rely heavily on hypothetical constructions and theoretical models. By shaping distributions with these theoretical methods, we can approximate the specific patterns that might emerge in reality. These hypothetical models are tools for current understanding and are invaluable in predicting future Photovoltaic (PV) penetration scenarios. By integrating forecasts and considering potential shifts in policy and technology, we can use these theoretical constructs to anticipate and prepare for changes in the energy landscape despite the current limitations in real-world data. It is worth noting, however, that the core objective of this paper diverges slightly. Rather than focusing on the details of creating these distributions, our primary interest lies in developing a framework to utilize these distributions to derive scenario-based insights on system reliability.
We define four different adoption types each for both PV and ES micro-level adoption. The adoption types are Limited (L), varied (V), median-focused (MF), and higly-concentrated (HC). Before discussing the choice of the distributions used to model the random variables, the range of the random variables needs to be defined. As of 2021, the average size of residential solar installations in the U.S. was approximately 7kW [21], compared to the average peak load of 4kW for a typical home, yielding a ratio of 1.75. This average ratio, indicative of a significant margin over immediate energy needs, also sets a reference for understanding the variability in solar energy production. By defining a range for X from 0 to 3.5, we accommodate the diversity of household energy profiles, acknowledging that some homes may not have solar capacity (hence the lower bound of 0). In contrast, others might have installations capable of generating up to twice the average output, providing a flexible scope that captures the spectrum of potential homeowner experiences.
Similarly, with residential energy storage, the popular residential energy storage solutions have capacities of 13.5kWh [21]. Considering this system as an average, it can support the peak load of an average household for approximately 3.375 hours. By extending the range from 0 to 6.75 hours, we encompass the full breadth of storage capacities, considering that some homes may opt not to have energy storage. In contrast, others may install systems that offer double the average duration of peak load coverage, thus ensuring preparedness for scenarios from short-term outages to extended periods without grid power.
These ranges are not arbitrary but thoughtfully chosen to represent the variability inherent in residential energy solutions. The average values give us a baseline, while the extended ranges ensure we capture the real-world fluctuations—both in terms of solar generation potential and energy storage capacity—thus providing a more comprehensive picture that homeowners can relate to and plan for.
The choice of distribution and the adoption type for both X and Y are described as follows and are visualized in Fig. 4.
Limited adoption: This adoption type is modeled using a scaled Beta distribution as follows.
$f(x^{\prime};\alpha,\beta)=\frac{1}{scale}\cdot\frac{x^{\prime\alpha-1}(1-x^{\prime})^{\beta-1}}{B(\alpha,\beta)}$ (39)
$x^{\prime}=\frac{x-loc}{scale}$ (40)
where $\alpha=0.75$ and $\beta=5$. The distribution is skewed towards the lower end of the range and the support is shifted from the standard [0,1] interval to [0, 3.5] for X and [0, 6.75] for Y by the $loc$ and $scale$ parameters.
  • X: This scenario assumes a conservative approach to the adoption of PV systems. Only a tiny fraction of homes have PV installations. Among those that do, the PV capacity is closer to the lower end of the defined range for $X$. In essence, both the number of houses with PV systems and the sizes of these systems are limited.
  • Y:Energy storage systems are not widely adopted. Among those with them, the storage capacity is minimal, possibly due to high costs, lack of awareness, or limited perceived need.
Varied adoption: This adoption type is modeled using a uniform distribution.
$f(x;a,b)=\begin{cases}\frac{1}{b-a}&\text{for }a\leq x\leq b\\ 0&\text{otherwise}\end{cases}$ (41)
where $a$, and $b$ are the range limits for the random variables.
  • X: Here, while the number of residential systems adopting PV might be more, the size of the installations varies greatly. Some houses might have a minimal PV capacity, while others approach the upper limit.
  • Y: Many homes have adopted energy storage systems, but their installed capacities vary greatly. While some homes might have minimal storage, others might approach the upper end of the capacity spectrum.
Figure 4: Marginal adoption scenarios for X.
Median-focused adoption: A truncated normal distribution is used to model the median-focused adoption.
$f(x;\mu,\sigma,a,b)=\frac{\phi\left(\frac{x-\mu}{\sigma}\right)}{\sigma\left(\Phi\left(\frac{b-\mu}{\sigma}\right)-\Phi\left(\frac{a-\mu}{\sigma}\right)\right)}$ (42)
for $a<x<b$, with $\phi$ and $\Phi$ denoting the PDF and the CDF of the standard normal distribution respectively. The middle value of the range ($\frac{a+b}{2}$) is taken as the mean ($\mu$) of the distribution with a relatively small standard deviation ($\sigma$). For X, $\sigma=0.5$ and for Y, $\sigma=1$.
  • X: The majority of homes with PV installations have capacities around a median value. This could be due to the standardization of the popular packages offered by installation companies.
  • Y: Most homes with energy storage systems have capacities close to the median value. This could be due to specific popular product offerings or household energy storage needs.
Highly concentrated: Beta distribution as defined in (39) is used to model this adoption scenario, where $\alpha=5$ and $\beta=0.75$. The shape parameters are modified to reflect high concentration near the upper bound of the range.
  • X: Here, most residential systems that adopt PV go for the higher end of the capacity spectrum. These are the homes maximizing their solar potential, possibly due to higher incentives, more significant energy needs, or other drivers.
  • Y: In this scenario, homes that adopt energy storage systems tend to install capacities on the higher end of the spectrum. Reasons could include larger homes, areas with unreliable grid power, or incentives that make larger systems more attractive.
V-C Joint adoption scenarios
TABLE II: Correlation for Joint Scenarios
X Y Limited Varied Median Highly Conc.
Limited 0.7 0.2 0.4 -0.2
Varied 0.2 0.4 0.3 0.1
Median 0.3 0.4 0.5 0.3
Highly Conc. 0.2 0.2 0.3 0.8
Given these marginal distributions, we construct joint adoption scenarios from these individual marginals. Table II provides a structured representation of hypothetical pairwise correlations between adoption scenarios for two variables, X and Y. Rows represents different adoption scenarios for Variable X. Columns represents different adoption scenarios for Variable Y. Values within the table indicate the assigned correlation coefficients between the corresponding scenarios of X and Y. It is essential to note that this table is not a correlation matrix. While a correlation matrix represents correlations between multiple variables, this table showcases the relationship between different scenarios of two specific variables. In essence, the table offers insights into how different adoption scenarios for X might relate to or influence corresponding scenarios for Y. It serves as a guide for understanding potential relationships between these variables under different circumstances.
V-D Application to RBTS Bus 4
The RBTS test system is made up of 7 feeders, 38 load points, and 4779 customers, of which 22 load points serve a total of 4700 residential customers [22]. The test system is modified as follows: Feeders 2,5, and 6, which serve nine industrial customers, are removed. Since the granularity of the simulation is at the first abstraction layer, the necessary condition is to have uniform reliability characteristics for all the load points across the distribution system. To achieve this, the system indices for the RBTS system, excluding the industrial feeders, are calculated. We use these indices as the basic load point indices for all the load points in the modified system. In this setup, we consider that commercial customers do not have any onsite DER. Therefore, the commercial customers’ AIF and AID will equal the load point failure and repair rate.
We chose the location of Los Angeles County for the simulation and calculated the representative residential load profile for the location from the ResStock dataset [23]. We used the solar irradiance data from [24], at the chosen location to calculate the rooftop PV system output in the simulation. We performed the MCS for the 16 joint adoption scenarios in Table II and tabulated the SAIFI and SAIDI values from the simulation results in Tables III and IV.
TABLE III: SAIFI for Joint scenarios (Baseline = 0.3 f/yr)
PV ES
Limited Varied Median Highly Conc.
Limited 0.2978 0.2982 0.29645 0.2962
Varied 0.2836 0.1894 0.2007 0.1317
Median 0.2850 0.1853 0.1759 0.0868
Highly Conc. 0.2738 0.1535 0.1463 0.0183
TABLE IV: SAIDI for Joint scenarios (Baseline = 3.47 hrs/yr)
PV ES
Limited Varied Median Highly Conc.
Limited 2.965 2.942 2.933 2.932
Varied 1.981 1.351 1.305 0.967
Median 1.884 1.172 0.954 0.511
Highly Conc. 1.729 0.923 0.703 0.095
Figure 5: Density plots on the left represent the Joint adoption density of PV and ES for the corresponding scenarios on the right. Density plots on the right represent the variation in AIF for the annotated scenarios.
The simulation results show the expected outcomes with SAIFI and SAIDI. As the adoption of PV and ES grows the system reliability metrics improve. Fig. 5, visualizes two selected adoption scenarios and the respective AIF index plots. The first set of subplots on the top correspond to the {Limited, Median-Focused} adoption scenario. The joint density contour on the left shows the PV and ES adoption concentration. The plot shows that the PV adoption is shallow between the range of having no PV to a PV size of one times the peak load. Meanwhile, the concentration of ES adoption is around the median value of 3. The corresponding density plot on the right shows the density of the AIF index constructed from the AIF values of the samples drawn from joint density. When we look at the estimated SAIFI value, we see a slight improvement from the baseline, but this does not tell the entire story. The variation in AIF, even though small, shows us that for some configurations of PV and ES sizes, the AIF deteriorates compared to the baseline. The following conditions explain the deterioration in the AIF values: if there is a long enough grid interruption, the end user with small PV and ES sizes will experience multiple interruptions due to insufficient power from the DER system and the cyclic nature of the PV generation. This phenomenon is reported in the literature from studies performed on a single residential unit. Our method quantifies and captures these intricate details at the residential level for an entire distribution system, which would otherwise be lost when a single metric is used to quantify system reliability.
The second set of subplots shows the {Varied, varied} adoption scenario. The joint density contours on the left show that the PV and ES adoption spans the entire range, and there is a high positive correlation. The corresponding density plot shows a wide variation in the AIF index. As with the previous scenario, the estimated AIF shows huge improvement, but the single numerical metric does not capture the entire variation in the AIF index. There are two peaks in the distribution, and the estimated SAIFI falls in between the peaks. The peak on the right of the estimated SAIFI has a higher density than the peak on the left, indicating that customers with higher PV and ES sizes contribute the most to the improvement in the estimated SAIFI. A single numerical metric quantifying the estimated SAIFI will not capture these variation and cannot provide detailed residential-level insights.While we simulated sixteen PV and ES joint scenarios, we explained only two scenarios due to space limitations. The rest of the scenarios are visualized in Figs. 5(a) and 5(b).
(a) SAIFI distributions for various adoption scenarios.
(b) SAIDI distributions for various adoption scenarios.
Figure 6: Each subplot is for a specific PV adoption and each distribution is for a specific ES adoption as indicated in the legend.
VI Conclusion
This work proposes a bottom-up probabilistic approach to integrate residential DER into reliability evaluation. The variability of the residential DERs is captured and integrated into the assessment using micro-level penetration modeling. Applying this approach to the RBTS test system shows that a single metric, such as SAIFI or SAIDI, will not capture the variation in reliability experienced by the end users in the presence of BTM DERs. The proposed approach provides a powerful way to quantify the variability in the reliability indices. The findings of this research are particularly relevant for utilities and policymakers, providing them with a robust framework to anticipate and manage the challenges posed by the rapid integration of residential DERs. Future work will further evaluate the impact of adopting this probabilistic approach in speeding up the transition towards more renewable energy resources in large-scale testing.